HOME

   


From the 1/14th Daily Journal Summaries

November 1968


1 November 1968

2055-2056: Cobalt Ripple Alpha is located at YA 703378. This information was passed to bde at 2243 (C).

2130-2131: To all elements from S-3. Effective 2400 tonight the Bn Command Frequency will change to the Command Bravo frequency (C).

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue to improve bunkers, fortifications and fields of fire on all LZís within out AO.

 

2 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue to improve bunkers, fortifications and fields of fire on all LZís within out AO.

 

3 November 1968

0315-0316: To Bn from Bde. SOI change-1-69 Armor command A frequency is 45.15 (C).

1025-1026: To Bn from Bde. Suffix change for 1-12 Inf from "I" to "L" (C).

2125-2126: To Bde from Bn. NL of Cobalt Ripple A is at 695339 (C).

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue SRP operations with fortifications of bunkers and construction of all LZís.

 

4 November 1968

0740-0741: To NCS from Bde. SOI change 1-69 Armor Company A 50.60 (C).

0930-0931: To S-3 from Commo Officer. C/1-14 new frequency is 43.95 (C).

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue SRP operations with fortifications of bunkers and construction of all LZís.

 

5 November 1968

1605-1606: To Bn from Bde. SOI changes. Bde LRRP Command A changes to 59.90 effective 05/1830. Command B to 62.30 effective 05/1900. Secure changed to 56.70 effective 05/1800. Bde will still be using Command A (C).

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue bunker construction and emphasis on fortifications of LZís. Also daily training will continue to be emphasized on all subjects.

 

6 November 1968

1130-1131: To NCS from S-2. Password for 061200 to 071200 is "ANNIE OVER".

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue emphasis will be placed on fortifications of bunkers and construction of the LZís. Also continue training emphasis for the individual units.

 

7 November 1968

2302-2303: To Bn from Bde. SOI change A/1-35 command A 66.25 effective 2115 7 Nov 68.

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue emphasis will be placed on fortifications of bunkers and construction of all LZís in our AO.

 

8 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue emphasis will be placed on fortifications of bunkers and construction of the LZís. An ambush is scheduled to be established at YA 784346 for the night of 9 Nov. On 10 Nov a bunker complex will be impregnated and partially destroyed with C/S powder.

 

9 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Continued preparations for the proposed operation to be conducted on 10 Nov 68. Emphasis was placed on the safe conduct of firing mortars.

Plans: Continue emphasis will be placed on fortifications of bunkers and construction of the LZís.

 

10 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Emphasis is still being placed on clearing fields of fire in all firebases.

Plans: Continue emphasis will be placed on fortifications of LZís and continued construction of all LZís.

 

11 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Additional emphasis is placed on SRP operations (route of march, noise and light disipline).

Plans: Continue construction and fortifications of LZís within their AO. Destruction plans for a bunker complex at YA 784346.

 

12 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Emphasis was also placed on SRP procedures along with ambushes.

Plans: Continue emphasis will be placed on fortifications of bunkers and construction of all LZís in the AO.

 

13 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Also a bunker complex was destroyed at YA 748346.

Plans: Continuing to send out SRPs and ambush patrols. Emphasis is being placed on overhead cover and blast walls on all bunkers in our AO.

 

14 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO.

Plans: Continue emphasis will be placed on fortifications of bunkers of all LZís. Also bunker construction will be increased and reinforced.

 

15 November 1968

Summary: The Golden Dragons continued SRP and ambush operations within their assigned AO.

Plans: B and C/1-14 will participate in BDA operations in the vicinity of YA7715. Improvement of fortifications continues at all LZís.

 

16 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. B/1-14 and C/1-14 departed LZ Karen to LZís located in vicinity of YA 778156 and moved to a night location. Movement was in conjunction with the units participation in BDA operations.

Plans: B and C/1-14 will continue BDA operations. SRP operations will be continued throughout assigned AO. Continued improvements of all Fire Bases.

 

17 November 1968

1545-1546: To Bn from Bde. Set of guns en route to cover your elements during move.

Summary: B and D/1-14 still engaged in BDA operations. The Golden Dragons continued SRP operations within their assigned AO.

Plans: B and D/1-14 will continue BDA operations. SRP operations will also be continued throughout the assigned AO.

 

18 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. B/1-14 and C/1-14 continue BDA operations.

Plans: B and C/1-14 will continue their participation in BDA operations. Continued emphasis on improvement of fortifications. SRP operations will continue throughout assigned AO.

 

19 November 1968

Summary: B/1-14 and C/1-14 returned to LZ Karen from BDA operations in which they had participated in since 16 Nov. SRP operations were also continued throughout the assigned AO.

Plans: The Golden Dragons will continue SRP operations throughout assigned AO. Improvements will continue on fortifications at all LZís.

 

20 November 1968

Summary: The Golden Dragons continued SRP operations within their assigned AO. Work was continued toward improvement of bunkers and fortifications and of the firebase in general.

Plans: Warning Order was received committing the Golden Dragons elements to participation in Operation "Hold Tight"; a cordon and search operation conducted by 3rd Bde. A and B/1-14 tentatively selected for this operation with support provided by elements from HHC. SRP operations will continue on a somewhat smaller scale due to our involvement in cordon and search operations.

 

21 November 1968

Summary: The Golden Dragons minus departed LZ Karen to participate in the Hold Tight, a cordon and search operation. Units participating are A and B/1-14. C/1-35(-). The LZ was secured without incident and no enemy contact was made through the day. SRP operations were cut back due to other commitments.

Plans: Improvement of bunkers and fortifications continues, SPR operations will continue in assigned AO.

 

22 November 1968

Summary: The Golden Dragons continued their participation in "Operation Hold Tight". At 1020 hours A/1-14 discovered a booby trapped area vicinity 047134. At 1215 A1-14 discovered another booby trapped area near a village vicinity 048136.

Plans: 1-14 will continue participation in "Operation Hold Tight". Units securing LZ Karen and Mary S will man the bunker lines stretching to cover areas normally manned by units absent through participation in current operations.

 

23 November 1968

Summary: The Golden Dragon elements participating in operation Hold Tight returned to their respective fire bases upon termination of that operation. B/1-14 returned to LZ Mary S, A/1-14 returned to LZ Karen. Reports of enemy activity within our assigned AO were as follows: #1-SRP #10 C/1-14 movement at 810303 (unconfirmed). #2-SRP #2 C/1014 movement at 025323. #3-C/1-14 movement 100 meters NW of LZ Mary S. #4-B/1-14 movement E of LZ Mary S. #5-SRP #8 C/1-14 camp fires at a distance of 125 meters on an azimuth of 270 degrees from his location at 826315. #6-SRP #4 C/1-14 movement on azimuth of 270 degrees at range of 500 meters. Organic overhead fire was employed in each case to reduce movement.

Plans: D/1-14 will organize a combat patrol of two combat platoons to patrol in the vicinity of 785326 where enemy activity was reported on 20 Nov. Improvements to firebase will continue with particular emphasis placed on bunkers and fortification of the LZís.

 

24 November 1968

Summary: The Golden Dragons continued SRP operations within assigned AO. General improvement to existing bunkers and fortifications continue and construction of several new bunkers was undertaken. Construction began on additional firing positions around the TOC.

Plans: A recon in force will be conducted during the period of 29 Nov to 1 Dec in the vicinity of LZ Karen.

 

25 November 1968

Summary: The Golden Dragons continued SRP and ambush operations within the assigned AO. D/1-14 conduced a recon in force to bunker complex, where enemy activity had previously been reported.

Plans: OP/LAN 30-68, HQ 3rd Bde directing the 1-14 Inf to participate in a search and destroy mission in Binh Tay-MacArthur west of LZ Vera and LZ Jean.

 

26 November 1968

Summary: The Golden Dragons continue preparations for the search and destroy operation in which selected elements of 1-14 are scheduled to participate. SRP operations continue within assigned AO. Incoming enemy mortar fire was received at LZ Jean at 0310. Other enemy activity reported is as follows: Recon SRP #3, enemy movement vicinity 793284. Improvement of fire base continues.

Plans: Plans are firm and required coordination has been made for 1-14 participation in search and destroy operation scheduled for tomorrow. B/1-14 and C/1-14 have been selected to take part, along with B/1-35 which is OP/CON to 1-14 for this operation. SRP operations will be reduced due to operational considerations.

 

27 November 1968

Summary: The Golden Dragons departed firebases for participation in operation STONEíS CRUSHER Ė a search and destroy mission. B/1-14 1st PZ at 0855. Last LZ 1045. C/1-14 1st PZ 0855 last LZ 0935. SRP operations continued within the assigned AO. A/1-14 SRP #8 had movement 50 to 75 meters SR of his location at 828303.

Plans: Continued emphasis will be placed on fortification of all LZís. SRP operations will stay at a minimum due to the fact that 2 companies area out on the operation.

 

28 November 1968

Summary: The Golden Dragons continued to participate in STONEíS CRUSHER, a search and destroy mission, involving 3rd Bde elements. Enemy contact was made by our B/1-14 in the afternoon and an enemy base camp was discovered by this same element. A helicopter carrying the Bn Co went down. Our rescue operations were promptly initiated and all passengers aboard were rescued with no casualties and without incidence. SRP operations continued within the assigned AO on a reduced scale and improvements to fire base continued.

Plans: Plans are being formulated for a 2 TF operation tentatively scheduled for period 30 Nov to 1 Dec in an AO east of LZ Karen. SRP operations will continue and improvement to fire base will continue.

 

29 November 1968

Summary: The Golden Dragons continue participation in Operation STONEíS CRUSHER. Every contact has been limited to sniper fire. Directives for relocation of selected 1-14 units have been received and relocation is in progress.

Plans: Relocation of units will continue. Orders have been received committing elements of 1-14 to participate in an operation north of the Oasis.

 

30 November 1968

Summary: The Golden Dragons continued field operations. B/1-14 and C/1-14 were extracted from STONEíS CRUSHER AO and air lifted to LZ in vicinity of Oasis to participate in operations north of that location. Relocation of elements of the 1-14 continues. The TOC will remain in place at LZ Karen.

Plans: Selected elements of 1-14 will continue their participation in operation north of Oasis. Restationing of units will continue.


TIMELINE 1968